# Sat-DN: Implicit Surface Reconstruction from Multi-View Satellite Images with Depth and Normal Supervision

### v1

@@ Abstract
With advancements in satellite imaging technology, acquiring high-resolution multi-view satellite imagery has become increasingly accessible, enabling rapid and location-independent ground model reconstruction. However, traditional stereo matching methods struggle to capture fine details, and while neural radiance fields (NeRFs) achieve high-quality reconstructions, their training time is prohibitively long. Moreover, challenges such as low visibility of building facades, illumination and style differences between pixels, and weakly textured regions in satellite imagery further make it hard to reconstruct reasonable terrain geometry and detailed building facades. To address these issues, we propose Sat-DN, a novel framework leveraging a progressively trained multi-resolution hash grid reconstruction architecture with explicit depth guidance and surface normal consistency constraints to enhance reconstruction quality. The multi-resolution hash grid accelerates training, while the progressive strategy incrementally increases the learning frequency, using coarse low-frequency geometry to guide the reconstruction of fine high-frequency details. The depth and normal constraints ensure a clear building outline and correct planar distribution. Extensive experiments on the DFC2019 dataset demonstrate that Sat-DN outperforms existing methods, achieving state-of-the-art results in both qualitative and quantitative evaluations. The code is available at https://github.com/costune/SatDN.
@@ I Introduction
We address these problems one by one. First, we propose to leverage monocular depth cues from a pre-trained depth estimation model[20] as an optimization constraint. We adjust the scale and offset of monocular relative depth using a point cloud refined by triangulation and bundle adjustment. Due to the smoothness of the predicted depth values, the disturbance of the illumination variation in the images on the depth values is minimal. We can use adjusted depth information to guide geometry, avoiding ambiguities in photometric optimization. Second, we fuse depth values to obtain monocular normal information and derive angular similarity from the normals. Non-edge pixels are treated as planar surfaces, and we apply a surface normal consistency constraint to guide the smooth distribution of normals, which helps achieve better results in large planar areas such as rooftops and roads. Third, considering that most buildings have regular appearances and fewer surface discontinuities, we introduce a multi-res hash feature grid[21] to store and optimize scene features, effectively reducing inference time. Additionally, we implement a progressive training strategy to prevent holes and indentations when reconstructing building facades.
@@
    Considering the common shapes of buildings, we integrate the multi-res hash feature grid into the pipeline. A progressive coarse-to-fine training strategy is also introduced to prevent the occurrence of holes and indentations in the reconstructed model.
@@
The subsequent part of the paper is structured as follows. Section II reviews the 3D reconstruction of satellite images and neural implicit representations. Section III details the preliminaries utilized in the proposed method. Section IV provides a comprehensive explanation of the structure and methodology adopted in Sat-DN. Section V presents the results of Sat-DN on the dataset and comparisons with other state-of-the-art (SOTA) methods. In section VI, we conclude the proposed method.
@@ III-A NeRF and NeuS
where $C_{k}$ and $\hat{C}_{k}$ denote the ground-truth color and rendered color, and the $\mathcal{L}_{eikonal}$ term is to regularize the SDF distribution.
@@ IV-A Depth Fusion and Regularization
Given a satellite image $I$ and a pre-trained depth model $F_{\theta}$, we can obtain the relative normalized dense depth of image $I$ as $\hat{D}_{dense}=F_{\theta}(I)$. The scale information is calculated from the sparse 3D points $\{P_{i}=(x_{i},y_{i},z_{i})\in\mathbb{R}^{3}\}_{i=1}^{M}$ obtained through image triangulation and bundle adjustment. Since the 3D points only contain information on longitude, latitude, and altitude, following [5], we obtain sparse depth by modeling the projection distance $D_{sparse}$ of the sparse points onto the reference plane where we reparameterize the ray’s start. Given a 3D sparse point $P$, whose projection on the 2D image is $p=(u,v)\in\mathbb{R}^{2}$, assuming the calculation process is modelled using the RPC model as:
$lon,lat=\text{RPC}(u,v,alt),$ (4)
where $lon,lat,alt$ means longitude latitude and altitude separately. We can derive the following:
$D_{sparse}=\left\|\left[\text{RPC}\right(p,alt\left),alt\right]-\left[\text{RPC}\left(p,alt_{ref}\right),alt_{ref}\right]\right\|_{2},$ (5)
where $alt$ means the altitude of the sparse point and $alt_{ref}$ means the altitude of the reference plane. The entire ray projection process is defined based on the UTM coordinate system, from the reference plane at the highest altitude to the reference plane at the lowest altitude. We treat the sparse points as the true terrain surface positions to correct the dense relative depth. Based on the dense relative depth and sparse absolute depth, we can obtain dense absolute depth as:
$D_{dense}=s\cdot\hat{D}_{dense}+o,$ (6)
where $s$, $o$ are scale value and offset value, computed through minimizing the differences between sparse absolute depth and fitted depth from the relative one:
$s^{*},o^{*}=\arg\min_{s,o}\sum_{i=1}^{M}|w(P_{i})\cdot D_{sparse}(P_{i})-D_{dense}(P_{i};s,o)|^{2},$ (7)
where $w\in[0,1]$ is a normalized weight representing the reprojection error of the point $P_{i}$ onto the image, namely, the fitting confidence for each point. We can use the dense absolute depth $D_{dense}=s^{*}\cdot\hat{D}_{dense}+o^{*}$ for pixel depth supervision after calculating $s^{*}$ and $o^{*}$. However, the depth model trained on a large set of standard photos is not always accurate when applied to satellite images. The depth predicted by the model for water surfaces creates discontinuities with surrounding land, causing depth supervision to be ineffective. Thus, we apply a semantic mask $M$ to exclude water regions with erroneous depth.
@@
$\mathcal{L}_{depth}=\frac{1}{K}\sum_{i=1}^{K}M_{i}|D_{i}^{pred}-D_{i}^{gt}|,$ (9)
where $D_{i}^{pred}$ is the predicted depth and $D_{i}^{gt}$ is the ground truth dense depth $D_{dense}$ for pixel $i$. $M_{i}\in\{0,1\}$ is the water mask. Similar to the pinhole camera, we transfer the altitude calculated from a RPC-based pushbroom camera into the depth with respect to the reference plane, which is suitable for normalization.
@@ IV-B Normal Regularization Based on Angular Consistency
$\delta_{i}^{gt}=\frac{1}{N}\sum_{j\in\text{Adj}(i)}\frac{\mathbf{n}_{i}\cdot\mathbf{n}_{j}}{\|\mathbf{n}_{i}\|\|\mathbf{n}_{j}\|},$ (10)
where $\text{Adj}(i)$ represents the neighbor pixels around the pixel $i$, $N$ is the total number of the neighbors.
@@
$\delta_{i}^{pred}=\frac{1}{N}\sum_{j\in\text{Adj}(i)}\frac{\nabla f(\mathbf{o}_{i}+t_{i}\mathbf{d}_{i})\cdot\nabla f(\mathbf{o}_{j}+t_{i}\mathbf{d}_{j})}{\|\nabla f(\mathbf{o}_{i}+t_{i}\mathbf{d}_{i})\|\|\nabla f(\mathbf{o}_{j}+t_{i}\mathbf{d}_{j})\|},$ (11)
where $\mathbf{o}$ and $\mathbf{d}$ means the origin and direction of the corresponding rays, $t_{i}$ is the rendered depth of the ray originated from pixel $i$. $N$ represents the number of adjacent pixels as above.
@@
$\mathcal{L}_{normal}=\frac{1}{K}\sum_{i=1}^{K}(\delta_{i}^{pred}-\delta_{i}^{gt})^{2},$ (12)
@@ IV-C Progressive Training Strategy
Due to the limited number of training images and constrained viewing angles, the surface model after training often contains holes and discontinuous artifacts beyond the surface contours. We observed that jointly optimizing low-resolution and high-resolution encodings tends to result in surface errors. This is because the model adjusts high-frequency features from the beginning of the train, making it prone to local optimization. Combined with the inherent issues of satellite imagery, this problem further harms accurate reconstruction. Therefore, we introduced a progressive training strategy, in which low-resolution grid features are activated first, and higher-resolution features are progressively activated as the training progresses. The feature representation for the sampled point $\mathbf{x}$ is defined as:
@@ V-A1 DFC2019
The DFC2019 (Data Fusion Contest 2019) dataset[22], is a multi-view satellite remote sensing dataset that includes hyperspectral imagery, multispectral imagery, LiDAR point cloud data, and other auxiliary data. The dataset contains satellite images captured over Jacksonville and Omaha, USA, by WorldView-3 over two years. Our task is to reconstruct the 3D surface model of the terrain based on the provided multi-view satellite imagery. The dataset includes DSMs generated from LiDAR point clouds, which serve as ground truth for reconstruction. Each DSM has dimensions of 512$\times$512 pixels with a resolution of 0.5 m, covering an area of 256m $\times$ 256m. To reduce the computational cost, we crop the original satellite images based on the DSM’s RoI and select five scenes from Jacksonville and Omaha for reconstruction. All image RPC parameters were refined through bundle adjustment. Additionally, we use supplementary data[59, 60] from the DFC2019 dataset to obtain semantic mask information for each image, which is also cropped according to the RoI. To ensure consistency in image style, we refrained from using winter-time satellite images from Omaha, avoiding the potential negative impact of style differences on the reconstruction results. The data parameters we used are shown in Table. I, and the scenes with their RoI are illustrated in Fig. 3.
@@
JAX_068 19 (30.347, -81.665) 17N 256$\times$256
JAX_175 24 (30.325, -81.638) 17N 256$\times$256
JAX_207 24 (30.315, -81.681) 17N 256$\times$256
JAX_214 24 (30.315, -81.665) 17N 256$\times$256
JAX_260 16 (30.311, -81.665) 17N 256$\times$256
OMA_203 25 (41.272, -95.952) 15N 256$\times$256
OMA_212 22 (41.276, -95.921) 15N 256$\times$256
OMA_248 23 (41.267, -95.934) 15N 256$\times$256
OMA_287 24 (41.253, -95.941) 15N 256$\times$256
OMA_315 23 (41.251, -95.855) 15N 256$\times$256
@@ V-B Implementation Details
[…]
@@ V-C3 S-NeRF and Sat-NeRF
V-C4 Sat-Mesh
@@
V-C5 FVMD-ISRe
FVMD-ISRe[10] effectively reconstructs geometry by improving the positional encoding strategy, incorporating the sunlight direction input into the color MLP, and integrating an RPC-based ray-casting approach into NeuS[12] reconstruction. We use the same number of input images as our method, with a batch size set to 4096 and 100K iterations, while other settings follow the default configuration.
@@ V-D Performance Comparison and Analysis
In Table. II, bold values in the table indicate the best results and the underlined values represent the second-best results. Arrows indicate the direction in which the evaluation metrics favor better results. In the JAX_068 and JAX_214 scenes, our method achieved the best MAE and CD results and was close to the best method in MED. In the JAX_175 scene, due to mismatches between the ground truth DSM and the real scene in the image, our method still produced the second-best results across all evaluation metrics, indicating that our method performs well in most regions of the scene with only slight discrepancies from the ground truth. In the JAX_207 and JAX_260 scenes, large vegetation and water surfaces affected the depth model’s depth estimation, leading to increased depth estimation errors that negatively impacted the reconstruction. However, our method still achieved relatively good results across all evaluation metrics.

### v2

@@ Abstract
With the advancement of satellite imaging technology, acquiring high-resolution multi-view satellite imagery has become increasingly accessible, enabling efficient and location-independent ground model reconstruction. However, traditional stereo matching methods often miss fine details, while neural radiance fields (NeRFs) provide high-quality reconstructions at the cost of prohibitively long training times. Moreover, challenges such as low visibility of building facades, illumination inconsistencies, and weakly textured regions further hinder accurate geometry recovery. To tackle these issues, we propose Sat-DN, a novel depth- and normal-guided surface reconstruction framework for satellite imagery. Given multi-view satellite images and predicted depth maps, Sat-DN employs a progressively trained multi-resolution hash grid architecture to efficiently represent scene geometry. Our pipeline integrates explicit depth supervision and surface normal consistency constraints to preserve structural boundaries and planar regularities. The progressive training strategy first fits low-frequency geometric structures and then refines high-frequency details under the guidance of depth and normals. Extensive experiments on the DFC2019 dataset demonstrate that Sat-DN significantly outperforms traditional MVS and NeRF-based baselines in both accuracy and efficiency. The code is publicly available at https://github.com/costune/SatDN.
@@ I Introduction
We address these problems one by one. First, we propose to leverage monocular depth cues from a pre-trained depth estimation model[20] as an optimization constraint. We adjust the scale and offset of monocular relative depth using a point cloud refined by triangulation and bundle adjustment. Due to the smoothness of the predicted depth values, the disturbance of the illumination variation in the images on the depth values is minimal. We can use adjusted depth information to guide geometry, avoiding ambiguities in photometric optimization. Second, we fuse depth values to obtain monocular normal information and derive angular similarity from the normals. Non-edge pixels are treated as planar surfaces, and we apply a surface normal consistency constraint to guide the smooth distribution of normals, which helps achieve better results in large planar areas such as rooftops and roads. Third, considering the smoothness inherent attribute of interpolation, we introduce a multi-res hash feature grid[21] to store and optimize scene features, effectively reducing inference time. Additionally, we employ a progressive training strategy to prevent holes and indentations when reconstructing building facades.
@@
    We integrate the multi-res hash feature grid into the pipeline to accelerate the training. A progressive coarse-to-fine training strategy is also introduced to prevent the occurrence of holes and indentations in the reconstructed model.
@@
The subsequent part of the paper is structured as follows. Section II reviews the 3D reconstruction of satellite images and neural implicit representations. Section III details the preliminaries utilized in the proposed method. Section IV provides a comprehensive explanation of the structure and methodology adopted in Sat-DN. Section V presents the results of Sat-DN on the dataset and comparisons with other state-of-the-art (SOTA) methods. In section VII, we conclude the proposed method.
@@ III-A NeRF and NeuS
where $C_{k}$ and $\hat{C}_{k}$ denote the ground-truth color and rendered color, $m$ and $n$ represent the number of rays and points in each ray, and the $\mathcal{L}_{eikonal}$ term is to regularize the SDF distribution.
@@ IV-A Depth Fusion and Regularization
Given a satellite image $I$ and a pre-trained depth model $F_{\theta}$, we can obtain the relative normalized dense depth of image $I$ as $\hat{D}^{\text{dense}}=F_{\theta}(I)$. The scale information is calculated from the sparse 3D points $\mathbf{P}=\{(lon_{i},lat_{i},alt_{i})\in\mathbb{R}^{3}|i=1,2,...N\}$ obtained through image triangulation and bundle adjustment. Since the 3D points contain only information on longitude, latitude, and altitude, following [5], we obtain sparse depth of a point $D^{\text{sparse}}_{i}$ by calculating the distance between itself and the start point of the ray that goes through the point from the re-parameterized reference plane. Given a 3D sparse point $P_{i}\in\mathbf{P}$, whose projection on the 2D image is $(u_{i},v_{i})\in\mathbb{R}^{2}$, assuming the image coordinate is derived from the corrected satellite RPC model. In addition, we can use RPC to calculate longitude and latitude reversely:
$lon_{i},lat_{i}=\text{RPC}(u_{i},v_{i},alt_{i}),$ (4)
where $lon_{i},lat_{i},alt_{i}$ means longitude, latitude and altitude of point $P_{i}$ separately. Since the projection coordinate $(u_{i},v_{i})$ does not align with the integer pixel, we approximate the projection $(u_{i}^{\prime},v_{i}^{\prime})=\text{round}(u_{i},v_{i})$. With the reprojection position $P_{i}^{\prime}=(lon_{i}^{\prime},lat_{i}^{\prime},alt_{i})=(\text{RPC}(u^{\prime}_{i},v^{\prime}_{i},alt_{i}),alt_{i})$ and ray’s start point $P_{i}^{\text{ref}}=(lon_{i}^{\text{ref}},lat_{i}^{\text{ref}},alt^{\text{ref}})=(\text{RPC}(u^{\prime}_{i},v^{\prime}_{i},alt^{\text{ref}}),alt^{\text{ref}})$, we can derive the following:
$D^{\text{sparse}}_{i}=\left\|P_{i}^{\prime}-P^{\text{ref}}_{i}\right\|_{2},$ (5)
where $alt^{\text{ref}}$ means the altitude of the upper reference plane. The entire ray projection process is defined based on the UTM coordinate system, from the reference plane at the highest altitude to the reference plane at the lowest altitude. We treat the sparse points as the true terrain surface positions to correct the dense relative depth. Based on the dense relative depth and sparse absolute depth, we can obtain dense absolute depth as:
$D^{\text{dense}}_{i}=s\cdot\hat{D}^{\text{dense}}_{i}+o,$ (6)
where $\hat{D}^{\text{dense}}_{i}$ is relative depth of pixel $(u_{i},v_{i})$ and $s$, $o$ are scale value and offset value, computed through minimizing the differences between sparse absolute depth and fitted depth from the relative one:
$s^{*},o^{*}=\arg\min_{s,o}\sum_{i=1}^{HW}|w_{i}\cdot D^{\text{sparse}}_{i}-D^{\text{dense}}_{i}(s,o)|^{2},$ (7)
where $w_{i}\in[0,1]$ is a normalized weight representing the reprojection error of the point $P_{i}$ onto the image, namely, the fitting confidence for each point. We can use the dense absolute depth $D^{\text{dense}}_{i}=s^{*}\cdot\hat{D}^{\text{dense}}_{i}+o^{*}$ for pixel depth supervision after calculating $s^{*}$ and $o^{*}$. However, the depth model trained on a large set of standard photos is not always accurate when applied to satellite images. The depth predicted by the model for water surfaces creates discontinuities with the surrounding land, making depth supervision ineffective. Thus, we apply a semantic mask $M$ to exclude water regions with erroneous depth in the training stage.
@@
$\mathcal{L}_{depth}=\frac{1}{K}\sum_{i=1}^{K}M_{i}|\tilde{D}_{i}-D^{\text{gt}}_{i}|,$ (9)
where $\tilde{D}_{i}$ is the predicted depth and $D_{i}^{\text{gt}}$ is the ground truth dense depth $D^{\text{dense}}$ for pixel $i$. $M_{i}\in\{0,1\}$ is the water mask and $K$ is the batch size. Analogous to an LLFF space, we transform the space between the upper and lower reference planes into a canonical space, which is suitable for optimization.
The depth in our method is reparameterized as the distance between a point on a ray and the ray’s origin on the upper reference plane. We fit the sparse reparameterized depth values from sparse points into a dense relative depth map predicted by the depth model. Due to the large distance between the satellite sensor and the Earth’s surface, and the limited perspective of the satellite, the variation between true depth and reparameterized depth is relatively small. This enables us to obtain an optimal least-squares solution through iterative refinement. To address issues caused by variations in pixel photometry, the depth is used to guide the training rather than to indicate the exact position of the ground-truth surface.
@@ IV-B Normal Regularization Based on Angular Consistency
$\delta_{i}^{\text{gt}}=\frac{1}{N}\sum_{j\in\text{Adj}(i)}\frac{\mathbf{n}_{i}\cdot\mathbf{n}_{j}}{\|\mathbf{n}_{i}\|\|\mathbf{n}_{j}\|},$ (10)
where $\text{Adj}(i)$ represents the neighbor 4 pixels around the pixel $i$, NN is the total number of the neighbors.
@@
$\delta_{i}^{\text{pred}}=\frac{1}{N}\sum_{j\in\text{Adj}(i)}\frac{\nabla f(\mathbf{o}_{i}+t_{i}\mathbf{d}_{i})\cdot\nabla f(\mathbf{o}_{j}+t_{i}\mathbf{d}_{j})}{\|\nabla f(\mathbf{o}_{i}+t_{i}\mathbf{d}_{i})\|\|\nabla f(\mathbf{o}_{j}+t_{i}\mathbf{d}_{j})\|},$ (11)
where $\mathbf{o}$ and $\mathbf{d}$ means the origin and direction of the corresponding rays, $t_{i}$ is the rendered depth of the ray originated from pixel $i$. NN represents the number of adjacent pixels as above.
@@
$\mathcal{L}_{normal}=\frac{1}{K}\sum_{i=1}^{K}(\delta_{i}^{\text{pred}}-\delta_{i}^{\text{gt}})^{2},$ (12)
@@ IV-C Progressive Training Strategy
Due to the limited number of training images and constrained viewing angles, the surface model after training often contains holes and discontinuous artifacts beyond the surface contours. We observed that jointly optimizing low-resolution and high-resolution encodings tends to result in surface errors. This is because the model adjusts high-frequency features from the beginning of the train, making it prone to local optimization. Moreover, in satellite imagery, the side facades of individual buildings appear only a limited number of times across all images and often exhibit color inconsistencies, resulting in a lack of consistent color supervision. Therefore, directly optimizing high-frequency details from the beginning is unstable and may lead to the fitting of a multimodal volume density field near the surface. In contrast, optimizing low-frequency features first allows the model to learn a roughly continuous surface distribution during the early stages of training, which helps to prevent significant surface deviations when high-frequency features are subsequently introduced. Therefore, we introduced a progressive training strategy, in which low-resolution grid features are activated first, and higher-resolution features are progressively activated as the training progresses. The feature representation for the sampled point $\mathbf{x}$ is defined as:
@@ V-A1 DFC2019
The DFC2019 (Data Fusion Contest 2019) dataset[22], is a multi-view satellite remote sensing dataset that includes hyperspectral imagery, multispectral imagery, LiDAR point cloud data, and other auxiliary data. The dataset contains satellite images captured over Jacksonville and Omaha, USA, by WorldView-3 over two years. Our task is to reconstruct the 3D surface model of the terrain based on the provided multi-view satellite imagery. The dataset includes DSMs generated from LiDAR point clouds, which serve as ground truth for reconstruction. Each DSM has dimensions of $\times512$ pixels with a resolution of 0.5 m, covering an area of 256m $\times$ 256m. To reduce the computational cost, we crop the original satellite images based on the DSM’s RoI and select five scenes from Jacksonville and Omaha for reconstruction. All image RPC parameters were refined through bundle adjustment. Additionally, we use supplementary data[59, 60] from the DFC2019 dataset to obtain semantic mask information for each image, which is also cropped according to the RoI. To ensure consistency in image style, we refrained from using winter-time satellite images from Omaha, avoiding the potential negative impact of style differences on the reconstruction results. The data parameters we used are shown in Table. I, and the scenes with their RoI are illustrated in Fig. 3.
@@
JAX_068 19 (30.347, -81.665) 17N $\times256$
JAX_175 24 (30.325, -81.638) 17N $\times256$
JAX_207 24 (30.315, -81.681) 17N $\times256$
JAX_214 24 (30.315, -81.665) 17N $\times256$
JAX_260 16 (30.311, -81.665) 17N $\times256$
OMA_203 25 (41.272, -95.952) 15N $\times256$
OMA_212 22 (41.276, -95.921) 15N $\times256$
OMA_248 23 (41.267, -95.934) 15N $\times256$
OMA_287 24 (41.253, -95.941) 15N $\times256$
OMA_315 23 (41.251, -95.855) 15N $\times256$
@@ V-B Implementation Details
For the hash grid[21], we use 24 levels with resolutions ranging from 16 to 2048, with a maximum hash table size of 219 and a feature dimension of 2 per level. The encoding function $\text{Emd}(.)$ applies a 26 position embedding and concatenates the normalized UTM space coordinates of the sampling points. The SDF MLP consists of 3 layers, taking the hashed features of the sampling point as input and outputting a 1-D SDF value and a 256-D intermediate feature. The color MLP also has 3 layers, with input as the sampling point coordinates, the gradient of the SDF value, the intermediate feature, and the 24 position-embedded ray direction and solar light direction. We follow NeuS[12] ray sampling strategy, sampling 128 points per ray. The experiment’s batch size is 4096, with 100K iterations, and all of the loss function weights $\lambda$ are set to 0.1. The normal neighbor number is 4. The output mesh is obtained by marching cubes, and the DSM is generated by the projection of the mesh. Depth-Anything-V2[20] is used in our method, which is trained on over 62.5M images with high generalization ability.
@@ V-C3 S-NeRF and Sat-NeRF
V-C4 SpS-NeRF
SpS-NeRF[62] derived a pixel-wise depth from MVS point cloud to regulate the sparse view optimization. We directly adopt the results reported in the paper.
V-C5 Sat-Mesh
@@
V-C6 FVMD-ISRe and NeuS
FVMD-ISRe[10] effectively reconstructs geometry by improving the positional encoding strategy, incorporating the sunlight direction input into the color MLP, and integrating an RPC-based ray-casting approach into NeuS[12] reconstruction. We use the same number of input images as our method, with a batch size set to 4096 and 100K iterations, while other settings follow the default configuration. The results of NeuS[12] follow the report on the FVMD-ISRe.
@@ V-D Performance Comparison and Analysis
In Table. II, bold values in the table indicate the best results and the underlined values represent the second-best results. Arrows indicate the direction in which the evaluation metrics favor better results. In the JAX_068 and JAX_214 scenes, our method achieved the best MAE and CD results and was close to the best method in MED. In the JAX_175 scene, due to mismatches between the ground truth DSM and the real scene in the image, our method still produced the second-best results across all evaluation metrics, indicating that our method performs well in most regions of the scene with only slight discrepancies from the ground truth. In the JAX_207 and JAX_260 scenes, large vegetation and water surfaces affected the depth model’s depth estimation, leading to increased depth estimation errors that negatively impacted the reconstruction. In the image, the surface of the water will appear local high reflection; Vegetation has no regular geometric structure, and its colors are relatively homogeneous and messy. This is not conducive to the depth model to estimate a reasonable relative depth. However, our method still achieved relatively good results across all evaluation metrics.